OpenStudy (anonymous):

9+43x(20-8) divided by 2 divided by 6 ((Btw the 43 is supposed to be 4 then exponet 3))

jhonyy9 (jhonyy9):

so 43 than will be 64 9+64*12/2/6 =9+64*6/6=9+64=73